# An averaging formula for Nielsen numbers of affine $n$valued maps on infra-nilmanifolds

### v1

@@ Introduction
Let $M$ be a closed manifold (or more generally a compact polyhedron) and let $f:M\to M$ be a self-map of $M$. In Nielsen-Reidemeister fixed point theory one attaches to each such map $f$ a number $N(f)$, the Nielsen number of $f$, which is a (sharp) lower bound for the number of fixed points of any map $g$ which is homotopic to $f$. In this paper we are concerned with infra-nilmanifolds. It is known that any self-map on an infra-nilmanifold is homotopic to a so-called affine map ([10]). These maps are rather algebraic in nature and one has developed an averaging formula for computing their Nielsen number ([8, 9]). Since the Nielsen number of a map is invariant under a homotopy, this formula allows us to compute the Nielsen number of any self-map on an infra-nilmanifold. For $n$-valued maps the situation is more complicated. It is no longer true that any $n$-valued map on an infra-nilmanifold is homotopic to an affine $n$-valued map ([6]), but nevertheless, these affine $n$-valued maps still form a large and interesting class of $n$-valued maps on infra-nilmanifolds. In this paper, we obtain a natural generalisation of the formula from [9] for computing the Nielsen number of any affine $n$-valued map on an infra-nilmanifold.
In the first section we recall the basic Nielsen theory for $n$-valued maps and describe a general averaging formula (1.1) for computing the Nielsen number of an $n$-valued map. In the second section, we recall the necessary concepts concerning infra-nilmanifolds and affine $n$-valued maps. In the third section we establish a decomposition of the fixed point set of an $n$-valued map on an infra-nilmanifold, which allows us to rewrite the general averaging formula into a more algebraic one (3.1). In the fourth section we further examine this formula in the case of affine $n$-valued maps, to finally arrive at the main result of this paper (Theorem 4.9), which is an averaging formula generalising in a natural way the formula from [9].
@@ 1. Nielsen theory of $n$-valued maps
$\varepsilon_{\alpha,i}=\left\{\begin{array}[]{ll}1&\text{if $p(\mathrm{Fix}(\alpha\tilde{f}_{i}))$ is an essential fixed point class of $f$,}\\ 0&\text{otherwise.}\end{array}\right.$
@@ 2. Infra-nilmanifolds and affine maps
The maps we consider in this paper are affine $n$-valued maps on infra-nilmanifolds. An infra-nilmanifold, as defined e.g. in [5], is the quotient of a connected and simply connected nilpotent Lie group $G$ under the action of an almost-Bieberbach group $\pi\subseteq G\rtimes\mathrm{Aut}(G)$, i.e. a torsion free cocompact discrete subgroup of $G\rtimes C$, with $C$ a maximal compact subgroup of $\mathrm{Aut}(G)$. Such an almost-Bieberbach group acts properly discontinuously on $G$ and hence $\pi$ is (isomorphic to) the fundamental group of the infra-nilmanifold $\pi\backslash G$. As before, we will view $\pi$ as the covering group of the universal covering $G\to\pi\backslash G$.
In the special case where this group $\pi$ is entirely contained in $G$, the quotient space is called a nilmanifold. Every infra-nilmanifold $\pi\backslash G$ is finitely covered by a nilmanifold $N\backslash G$: it follows from the generalised Bieberbach Theorems that $\pi\cap G$ is of finite index in $\pi$, so one can take $N=\pi\cap G$ or any other subgroup $N\subseteq\pi\cap G$ which is of finite index in $\pi$. Such a group $N$ is a finitely generated torsion free nilpotent group and is a lattice of the Lie group $G$.
@@
Let $N\backslash G$ be a nilmanifold. The Nielsen number of an affine map $f:N\backslash G\to D_{n}(N\backslash G)$ with lift $\tilde{f}:G\to F_{n}(G,N):x\mapsto(g_{1}\varphi_{1}(x),\ldots,g_{n}\varphi_{n}(x))$ is given by
@@
The goal of this paper is to generalise this formula to infra-nilmanifolds. In the single-valued case, this extension from nilmanifolds to infra-nilmanifolds was done by Kim, Lee and Lee in [8]. They proved that any map $f$ on an infra-nilmanifold $\pi\backslash G$ can be lifted to a nilmanifold $N\backslash G$ that finitely covers $\pi\backslash G$, and they expressed the Nielsen number of $f$ as the average of the Nielsen numbers of these lifts.
@@ 3. Decomposition of the fixed point set
Let $\pi\backslash G$ be an infra-nilmanifold, finitely covered by a nilmanifold $N\backslash G$. For a general map $f:\pi\backslash G\to D_{n}(\pi\backslash G)$, we will decompose the sum (1.1) into a sum over Reidemeister classes in $N$ and in $\pi/N$, generalising the approach in [8].
For all $i$, consider the group $S^{\prime}_{i}$ generated by $\{\gamma^{[\pi:N]}\mid\gamma\in S_{i}\}$. This is a normal subgroup of $S_{i}$, and the quotient group $S_{i}/S^{\prime}_{i}$ is periodic, since the order of every element divides $[\pi:N]$. Also note that $\pi$ (and therefore also $S_{i}/S^{\prime}_{i}$) is polycyclic-by-finite, since its finite index normal subgroup $N$ is finitely generated and nilpotent. Since periodic polycyclic-by-finite groups are automatically finite, it follows that $[S_{i}:S^{\prime}_{i}]$ is finite. Consequently, also
$[N:S^{\prime}_{i}]=\frac{[\pi:S_{i}][S_{i}:S^{\prime}_{i}]}{[\pi:N]}$
@@
If $[g\alpha]=[g^{\prime}\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$, then certainly $[\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, i.e.
@@
Conversely, note that if $[\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, then there is a $g\in N$ so that $[g\alpha]=[\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$.
For fixed $\alpha$, we have $[g\alpha]=[g^{\prime}\alpha]$ in $\mathcal{R}[\phi_{i}]$ if and only if $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, with $\tau_{\alpha}:\pi\to\pi:\beta\mapsto\alpha\beta\alpha^{-1}$; i.e.
@@
Since $S_{i}^{\prime}\subseteq S_{i}$, it is clear that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ implies $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, but conversely there can be distinct elements in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ that are equivalent in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. The following lemma allows us to describe how many different classes in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ give rise to the same class in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. In this lemma, we use $\mathrm{coin}(\varphi,\psi)=\{\gamma\in\Pi_{1}\mid\varphi(\gamma)=\psi(\gamma)\}$ to denote the group of coincidence points of two morphisms $\varphi,\psi:\Pi_{1}\to\Pi_{2}$.
@@
such that the quotient groups $\Pi_{i}/\Gamma_{i}$ are finite. For all $\alpha\in\Pi_{2}$ and $g\in\Gamma_{2}$, the number of Reidemeister coincidence classes $[g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\varphi^{\prime},\psi^{\prime}]$ such that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\varphi,\psi]$ is equal to
$[\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\varphi},\bar{\psi}):\overline{\mathrm{coin}(\tau_{g\alpha}\varphi,\psi)}]$
@@
learns that for a given $g\in N$, the number of classes $[g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ such that $[g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$ is equal to
$[\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota}):\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}]=\frac{|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|}{|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|},$
@@
The orbit of an element $\bar{\alpha}$ for this action is its Reidemeister class $[\bar{\alpha}]\in\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$. The stabilizer of $\bar{\alpha}$ is the set
@@ 4. Averaging formula for affine maps
Suppose $\pi\backslash G$ is an infra-nilmanifold finitely covered by a nilmanifold $N\backslash G$, and $S$ is a finite index subgroup of $\pi$. Let $S^{\prime}$ be the subgroup generated by $\{\gamma^{[\pi:N]}\mid\gamma\in S\}$. Let $\phi:S\to\pi$ be a morphism and $\phi^{\prime}:S^{\prime}\to N$ its restriction to $S^{\prime}$. If $\det(I-(\phi^{\prime})_{*})\neq 0$, then $\mathrm{fix}(\phi)=1$.
@@
Suppose $\gamma\in S^{\prime}$ is a fixed point of $\phi^{\prime}$. Then $X=\log(\gamma)$ is a fixed point of $(\phi^{\prime})_{*}$, i.e. $(I-(\phi^{\prime})_{*})(X)=0$. Since $\det(I-(\phi^{\prime})_{*})\neq 0$, the only element $X$ satisfying this equation is 00. It follows that $\gamma=\exp(0)=1$.
@@
In [4, Lemma 6.10] it is proven that, if $\varphi:G\to G$ is a morphism so that $G\to G:x\mapsto\varphi(x)x^{-1}$ is a bijection, then for any $h\in G$, the index of $G\to G:x\mapsto h\varphi(x)$ at its unique fixed point is ±$\pm 1$. Applying this to $\varphi=A_{\alpha}\varphi_{i}$ and $h=g^{\prime}$ gives the desired result. ∎
@@
$\lambda v_{1}+\mu v_{1}+(1-\lambda)v_{1}=\mu v_{1}\;\mathrm{mod}\,V,$
@@
$\pi\backslash G\times[0,1]\to D_{n}(\pi\backslash G):(p(x),t)\mapsto\{p(\tilde{f}_{1}(x)g_{0}^{t}),\ldots,p(\tilde{f}_{n}(x)g_{0}^{t})\}.$
Under this homotopy, the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ of $f$ corresponds to the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}g_{0}))$. Since $g\alpha\tilde{f}_{i}g_{0}$ has no fixed points by definition of $g_{0}$, the index of this fixed point class is 00. As indices are preserved under homotopies, the index of $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ is also 00. ∎
@@ 5. Example
$a=\left(\begin{bmatrix}1\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right),\quad b=\left(\begin{bmatrix}0\\ \frac{1}{2}\end{bmatrix},\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\right).$
@@
This map induces a well-defined $2$-valued map $f:\pi\backslash\mathbb{R}^{2}\to D_{2}(\pi\backslash\mathbb{R}^{2})$ on the Klein bottle, since for all $(t_{1},t_{2})\in\mathbb{R}^{2}$
@@
with $\sigma\in\Sigma_{2}$ the permutation that interchanges $1$ and $2$.
@@
To compute the Nielsen number of $f$, we consider the nilmanifold $Z\backslash\mathbb{R}^{2}$ where $Z=\pi\cap\mathbb{R}^{2}=\langle a,b^{2}\rangle\cong\mathbb{Z}^{2}$, i.e. the $2$-torus. The group $\pi/Z$ has two elements, represented by
$\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right)\quad\text{and}\quad\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\right).$
@@
$N(f)$ $=\frac{1}{2}\sum_{\bar{\alpha}\in\pi/Z}\sum_{i=1}^{2}\left|\,\det\left(I-A_{\alpha}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$
$=\sum_{\bar{\alpha}\in\pi/Z}\left|\,\det\left(I-A_{\alpha}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$
$=\left|\,\det\left(I-\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|+\left|\,\det\left(I-\begin{bmatrix}-1&0\\ 0&1\end{bmatrix}\begin{bmatrix}0&0\\ 0&\frac{1}{2}\end{bmatrix}\right)\right|$

### v2

# An averaging formula for Nielsen numbers of affine $n$ -valued maps on infra-nilmanifolds

@@ Introduction
Let $M$ be a closed manifold (or more generally a compact polyhedron) and let $f:M\to M$ be a self-map of $M$. In Nielsen-Reidemeister fixed point theory one attaches to each such map $f$ a number $N(f)$, the Nielsen number of $f$, which is a (sharp) lower bound for the number of fixed points of any map $g$ which is homotopic to $f$. In this paper we are concerned with infra-nilmanifolds. It is known that any self-map on an infra-nilmanifold is homotopic to a so-called affine map ([10]). These maps are rather algebraic in nature and one has developed an averaging formula for computing their Nielsen number ([8, 9]). Since the Nielsen number of a map is invariant under a homotopy, this formula allows us to compute the Nielsen number of any self-map on an infra-nilmanifold. For $n$-valued maps we also have the concept of affine maps (see [6] and/or section 2), but there the situation is more complicated. It is no longer true that any $n$-valued map on an infra-nilmanifold is homotopic to an affine $n$-valued map ([6]), but nevertheless, these affine $n$-valued maps still form a large and interesting class of $n$-valued maps on infra-nilmanifolds. In this paper, we obtain a natural generalisation of the formula from [9] for computing the Nielsen number of any affine $n$-valued map on an infra-nilmanifold.
In the first section we recall the basic Nielsen theory for $n$-valued maps and describe a general averaging formula (1.1) for computing the Nielsen number of an $n$-valued map. In the second section, we recall the necessary concepts concerning infra-nilmanifolds and affine $n$-valued maps. In the third section we establish a decomposition of the fixed point set of an $n$-valued map on an infra-nilmanifold, which allows us to rewrite the general averaging formula into a more algebraic one (3.1). In the fourth section we further examine this formula in the case of affine $n$-valued maps, to finally arrive at the main result of this paper (Theorem 4.9), which is an averaging formula generalising in a natural way the formula from [9]. Finally in the last section we illustrate our main result by means of an affine 2-valued map on the Klein bottle. Moreover, the example we provide is special in the sense that this map does not lift to a 2-valued map on any torus (nilmanifold) that finitely covers the Klein bottle.
@@ 1. Nielsen theory of $n$-valued maps
[…]
@@ 2. Infra-nilmanifolds and affine maps
The maps we consider in this paper are affine $n$-valued maps on infra-nilmanifolds. An infra-nilmanifold, as defined e.g. in [5], is the quotient of a connected and simply connected nilpotent Lie group $G$ under the action of an almost-Bieberbach group $\pi\subseteq G\rtimes\mathrm{Aut}(G)$, i.e. a torsion free cocompact discrete subgroup of $G\rtimes C$, with $C$ a maximal compact subgroup of $\mathrm{Aut}(G)$. Such an almost-Bieberbach group acts properly discontinuously on $G$ and hence $\pi$ is (isomorphic to) the fundamental group of the infra-nilmanifold $\G\pi\backslash$ G. As before, we will view $\pi$ as the covering group of the universal covering $\GG\to\pi\backslash G$.
[…]
@@
[…]
@@
[…]
@@ 3. Decomposition of the fixed point set
[…]
For all $i$, consider the group $S^{\prime}_{i}$ generated by $\{\gamma^{[\pi:N]}\mid\gamma\in S_{i}\}$. This is a normal subgroup of $S_{i}$, and the quotient group $S_{i}/S^{\prime}_{i}$ is periodic, since the order of every element divides $\pi:N]$. Also note that $\pi$ (and therefore also $S_{i}/S^{\prime}_{i}$) is polycyclic-by-finite, since its finite index normal subgroup $N$ is finitely generated and nilpotent. Since periodic polycyclic-by-finite groups are automatically finite, it follows that $S_{i}:S^{\prime}_{i}]$ is finite. Consequently, also
$N:S^{\prime}_{i}]=\frac{[\pi:S_{i}][S_{i}:S^{\prime}_{i}]}{[\pi:N]}$
@@
If $g\alpha]=[g^{\prime}\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$, then certainly $\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, i.e.
@@
Conversely, note that if $\bar{\alpha}]=[\bar{\alpha}^{\prime}]$ in $\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$, then there is a $g\in N$ so that $g\alpha]=[\alpha^{\prime}]$ in $\mathcal{R}[\phi_{i}]$.
For fixed $\alpha$, we have $g\alpha]=[g^{\prime}\alpha]$ in $\mathcal{R}[\phi_{i}]$ if and only if $g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, with $\tau_{\alpha}:\pi\to\pi:\beta\mapsto\alpha\beta\alpha^{-1}$; i.e.
@@
Since $S_{i}^{\prime}\subseteq S_{i}$, it is clear that $g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ implies $g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$, but conversely there can be distinct elements in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ that are equivalent in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. The following lemma allows us to describe how many different classes in $\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ give rise to the same class in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$. In this lemma, we use $\mathrm{coin}(\varphi,\psi)=\{\gamma\in\Pi_{1}\mid\varphi(\gamma)=\psi(\gamma)\}$ to denote the group of coincidence points of two morphisms $\varphi,\psi:\Pi_{1}\to\Pi_{2}$.
@@
such that the quotient groups $\Pi_{i}/\Gamma_{i}$ are finite. For all $\alpha\in\Pi_{2}$ and $g\in\Gamma_{2}$, the number of Reidemeister coincidence classes $g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\varphi^{\prime},\psi^{\prime}]$ such that $g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\varphi,\psi]$ is equal to
$\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\varphi},\bar{\psi}):\overline{\mathrm{coin}(\tau_{g\alpha}\varphi,\psi)}]$
@@
learns that for a given $g\in N$, the number of classes $g^{\prime}]\in\mathcal{R}[\tau_{\alpha}\phi^{\prime}_{i}]$ such that $g]=[g^{\prime}]$ in $\mathcal{R}[\tau_{\alpha}\phi_{i}]$ is equal to
$\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota}):\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}]=\frac{|\mathrm{coin}(\tau_{\bar{\alpha}}\bar{\phi}_{i},\bar{\iota})|}{|\overline{\mathrm{fix}(\tau_{g\alpha}\phi_{i})}|},$
@@
The orbit of an element $\bar{\alpha}$ for this action is its Reidemeister class $\bar{\alpha}]\in\mathcal{R}[\bar{\phi}_{i},\bar{\iota}]$. The stabilizer of $\bar{\alpha}$ is the set
@@ 4. Averaging formula for affine maps
[…]
@@
Suppose $\gamma\in S^{\prime}$ is a fixed point of $\phi^{\prime}$. Then $X=\log(\gamma)$ is a fixed point of $(\phi^{\prime})_{*}$, i.e. $(I-(\phi^{\prime})_{*})(X)=0$. Since $\det(I-(\phi^{\prime})_{*})\neq 0$, the only element $X$ satisfying this equation is 0. It follows that $\gamma=\exp(0)=1$.
@@
In [4, Lemma 6.10] it is proven that, if $\varphi:G\to G$ is a morphism so that $G\to G:x\mapsto\varphi(x)x^{-1}$ is a bijection, then for any $h\in G$, the index of $G\to G:x\mapsto h\varphi(x)$ at its unique fixed point is $\pm 1$. Applying this to $\varphi=A_{\alpha}\varphi_{i}$ and $h=g^{\prime}$ gives the desired result. ∎
@@
$\lambda v_{1}+\mu v_{1}+(1-\lambda)v_{1}\equiv\mu v_{1}\;\mathrm{mod}\,V,$
@@
$\pi\backslash G\times[0,1]\to D_{n}(\pi\backslash G):(p(x),t)\mapsto\{p(g\alpha\tilde{f}_{1}(x)g_{0}^{t}),\ldots,p(g\alpha\tilde{f}_{n}(x)g_{0}^{t})\}.$
Under this homotopy, the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ of $f$ corresponds to the fixed point class $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}g_{0}))$. Since $g\alpha\tilde{f}_{i}g_{0}$ has no fixed points by definition of $g_{0}$, the index of this fixed point class is 0. As indices are preserved under homotopies, the index of $p(\mathrm{Fix}(g\alpha\tilde{f}_{i}))$ is also 0. ∎
@@ 5. Example
[…]
@@
This map induces a well-defined 22-valued map $f:\pi\backslash\mathbb{R}^{2}\to D_{2}(\pi\backslash\mathbb{R}^{2})$ on the Klein bottle, since for all $(t_{1},t_{2})\in\mathbb{R}^{2}$
@@
with $\sigma\in\Sigma_{2}$ the permutation that interchanges 11 and $2$.
@@
To compute the Nielsen number of $f$, we consider the nilmanifold $Z\backslash\mathbb{R}^{2}$ where $Z=\pi\cap\mathbb{R}^{2}=\langle a,b^{2}\rangle\cong\mathbb{Z}^{2}$, i.e. the 22-torus. The group $\pi/Z$ has two elements, represented by
[…]
@@
[…]